OpenStudy (jenniferjuice):

- 1/2 x + 2 = −x + 7

OpenStudy (whpalmer4):

\[-\frac{1}{2}x + 2 = -x + 7\]Is that correct?

OpenStudy (whpalmer4):

Let's assume it is. When faced with a problem that has fractions in it, I usually multiply the whole thing by the least common multiple of all of the denominators to make the fractions go away. Here we only have the 1 fraction, so multiplying by its denominator will suffice.